# Inequality in Congestion Games with Learning Agents

### v1

@@ Abstract.
Key words and phrases:
Congestion Games, Reinforcement Learning, Multi-Agent Systems, Fairness, Transport Planning
1. Introduction
Expanding transport networks is often seen as an unambiguous good: more roads, more transport lines, more options means better access and faster travel. Yet, in practice, such interventions can backfire. Sometimes, adding infrastructure slows traffic—as showcased in Braess’s Paradox braess_paradox—and other times, it disproportionately benefits already advantaged areas, widening inequalities martens2016transport.
@@ 1. Introduction
These challenges expose a critical blind spot in how we model and evaluate urban mobility. Classical game theoretic models assume homogeneous, rational agents with perfect information. But real commuters behave differently: they experiment and adapt. And although research that investigates this type of adaptation are increasing carissimo_counter_intuitive; selten_commuters_2007; ben-elia_which_2010, they have not yet been analyzed through the lens of fairness, especially when agents differ in their capacity to learn. In this paper, we use multi-agent reinforcement learning on congestion games to investigate how exploration-driven learning shapes the distributional outcomes of transport expansions.
@@
A well-known example of unintended outcomes is the Braess’s Paradox, which shows that adding a new route can paradoxically worsen congestion. Initially a theoretical result, the paradox has since been observed in real-world cities like New York, Boston, and Seoul youn2008price; baker2009removing. Such paradoxes illustrate the power of game-theoretic models, which capture mobility as a strategic interaction among selfish agents roughgarden_severity_2006; Zhang_coordinated_route_2023.
@@
Yet these models rest on the assumption of perfectly rational, fully informed commuters. In practice, travellers adapt through trial and error: experimenting with routes, observing outcomes, and adjusting their choices over time. This process closely resembles Reinforcement Learning (RL), in which agents balance exploration (trying new strategies) with exploitation (relying on known ones) sutton_reinforcement_2018. Empirical studies have shown that even simple RL mechanisms can successfully reproduce observed patterns in commuter behaviour selten_commuters_2007; ben-elia_which_2010. Furthermore, modeling commuter behaviour through learning dynamics has shown deviations from the expected outcomes of full-information strategic agents — agents do not necessarily converge to the Nash Equilibrium wunder2010classes; carissimo_counter_intuitive. As such, introducing learning dynamics into traffic models not only increases realism but can also reshape expectations about system performance. Moreover, it enables the analysis throughout the entire adaptive process, rather than restricting attention solely to equilibria. We extend this line of research by tackling the fundamental question: How do differences in learning affect fairness and efficiency in congestion games?
@@
Our question is further motivated by empirical research suggesting that wealthier individuals showcase a higher risk tolerance, and are generally more exploratory in their behaviours dong2020segregated; falleur_income_behavior. For instance, dong2020segregated find that residents of low-income neighborhoods are significantly less exploratory in their activities, effectively living within constrained environments. Psychological studies indicate that individuals from high-resource background demonstrate a higher tolerance for uncertainty and risk falleur_income_behavior. These findings imply that exploration and learning rates can act as a proxies for resource capacity: individuals with greater resources are better equipped to search for and evaluate alternatives, while those with fewer resources are constrained by limitations on time, mobility, and finances. Such heterogeneity can exacerbate inequalities, yet it remains underexplored.
@@ 2. Related Work
2.1. Reinforcement Learning in Congestion Games
@@
Deviations from the standard rational commuter behaviour have been studied under models such as prospect theory, which consider uncertainty in travelers’ decision-making chremos2022design; chremos_traveler-centric_2023. Recent papers revealed non-Nash convergence and even chaotic behaviour. For instance, online learning algorithms such as Multiplicative Weights Update chotibut2020route and Follow-the-Regularized-Leader bielawski2021follow demonstrate that deviations from Nash equilibrium occur when considering learning agents. Learning automata have also been used de_o_ramos_improved_2015.
@@
Reinforcement Learning is increasingly used to model travel choice behaviour in multiple congestion games ahmad2024travel. Examples include studies on how commuters adapt over time through learning dynamics cats2020learning; ben-elia_which_2010, the impact of memory wei_day_to_day, as well as the impact of tolls in congestion games ramos2020toll.
@@
Analyzing the effects of learning in congestion games through Reinforcement Learning, and in particular, multi-agent Q-learning hasn’t been explored thoroughly. Some efforts have achieved this in a parallel road network, showing that the total social cost varies with different exploration rates levy2018emergence. In particular, they show that with sufficiently small exploration, the system will converge to the inefficient equilibrium. Others investigate the effect of learning on a system’s convergence between Nash equilibrium and the social optimum carissimo_counter_intuitive. However, while they assume a homogeneous population with a uniform learning rates, in reality, commuters can show heterogeneity in learning. Here, we build upon this work and extend this by modeling two distinct populations with varying learning rates, allowing for a richer analysis of strategic behaviour and its implications for equilibrium outcomes, on both a stylilzed and a real-world network.
2.2. Fairness in Congestion Games
@@
The challenge of avoiding increased costs after network expansions, as showcased by the well-known Braess’s Paradox, has been extensively studied assuming selfish agents, with efforts focused on network design roughgarden_severity_2006; gairing_network_design_2017; roman2019multi.
@@
Our work contributes to this literature by examining fairness considerations in congestion games, an aspect previously explored primarily in centrally coordinated settings. Prior research has explored fairness through various mechanisms. One line of work analyzed the use of taxation to balance fairness and efficiency ferguson2021impact, while another investigated artificial currency–based pricing schemes as a means to promote equitable outcomes pedroso2024fair. A karma-based framework for achieving fairness was proposed in related studies censi2019today. Further, the Double Braess’s Paradox was introduced in a multi-resource setting with two sources and two destinations, where action-restriction mechanisms were evaluated as a fairness intervention oesterle2024raise. Additional work on weighted congestion games has examined mechanisms designed to reduce disparities in agent outcomes fischer2023fair.
@@
Using agent-based models, prior work has examined how individual behaviors affect fairness and efficiency in network settings. One study analyzed a five-edge Braess’s paradox network, evaluating the impact of various micro-level parameters on system performance Belov04072022. Another line of research investigated fairness in travel times between two routes through the lens of learning dynamics, assuming homogeneous exploration and learning rates levy2018emergence.
@@
3. Preliminaries
3.1. Congestion Games
We consider standard congestion games roughgarden_severity_2006. Let $G=(V,E)$ be a directed graph, where $V$ is the set of nodes and $E$ the set of edges. Let $\mathcal{N}$ be the set of players with $N=|\mathcal{N}|$, where each player $i\in\mathcal{N}$ has a source node $s_{i}\in V$ and a common destination $d\in V$.
@@ 3.1. Congestion Games
  • Continuity and Monotonicity: $f_{e}$ is continuous and nondecreasing roughgarden2002unfair.
  • Polynomial-time computation: $f_{e}(x)$ can be computed in polynomial time fischer2023fair.
@@ 3.2. Nash Equilibrium
Since congestion games are potential games monderer_potential_1996; milchtaich1996congestion, at least one pure-strategy NE exists rosenthal1973network.
3.3. Social Optimum, Price of Anarchy, and Price of Learning
@@
The Price of Anarchy (PoA) measures inefficiency of NE relative to the social optimum roughgarden_2002_how_bad_selfish_routing:
@@
where $a^{\star}$ is the socially optimal profile (Equation 6).
@@
3.4. Fairness
We assess fairness through the notion of source disparity, defined as the difference in average cost between groups of players originating from different sources. Our definition is inspired by fairness (or justice) concepts in urban planning, where location is central: geographic position determines accessibility to amenities and thus shapes opportunities martens2016transport. Our metric requires that no source node is systematically disadvantaged in the costs incurred to reach a destination. From the perspective of artificial intelligence fairness, this corresponds to the group fairness criterion of independence, which demands that outcomes—here, travel costs—remain independent of an agent’s point of origin castelnovo_clarification_2022.
@@ 3.4. Fairness
4. Learning Dynamics
In practice, humans deviate from exact Nash equilibrium strategies, as they lack complete information about which strategies yield specific payoffs bonau_begavioural_game_theory; lee_game_2008. Instead of playing optimally from the outset, individuals tend to explore various options and gradually learn to favor those that maximize their satisfaction. This adaptive behavior has been studied in congestion games, where outcomes frequently diverge from Nash equilibrium predictions carissimo_counter_intuitive. For this reason, we argue that it is essential to study the process itself, rather than assuming that commuters will ultimately converge. To achieve this, we use a framework in which agents repeatedly play the game and adapt their strategies through Reinforcement Learning (RL).
Importantly, we employ RL not as a tool for optimisation, but to simulate the adaptive behaviours of agents interacting within complex environments. Our perspective aligns with the descriptive agenda in multi-agent learning, which emphasizes modeling learning processes that mirror how real-world agents, such as humans, adapt in social contexts SHOHAM2007365. Importantly, this focus is not limited to human learning. The same framework extends to RL-based systems themselves. For example, as RL-powered recommender systems become increasingly prevalent afsar_rl_survey, it is likely that real-time adaptive systems will also be deployed in navigation applications. Understanding their learning dynamics is therefore equally crucial.
@@ 4.1. Q-Learning Framework
with $a_{i}(t)$ the strategy chosen at time $t$ and $a_{t}$ the realized profile. The negative sign ensures Q-learning maximizes the negative latency (minimizes actual latency). We assume the game is stateless and rewards are observed immediately (as in, e.g., claus1998dynamics; carissimo_counter_intuitive). Agents starting in different sources may have different learning rates.
@@
Convergence At each step, agents collectively contribute to congestion. Each agent estimates the expected cost of selecting a given strategy. With continued learning, Q-values converge to estimates reflecting expected cost. However, these do not necessarily correspond to a Nash Equilibrium (NE); convergence is shaped by factors such as the learning rate, exploration rate, and the environment’s structural properties carissimo_counter_intuitive. Unlike traditional RL aimed at maximizing a reward, our focus here is on describing and understanding emergent phenomena in a population of learning agents.
Figure 1. A Braess’s Paradox game with two sources ($S1$ & $S2$) connecting to a single destination node $B$. We examine two phases: before and after the network extension $I_{CD}$ (red link).
5. Environments
We introduce two routing games: a stylized two-source extension of Braess’s paradox (Section 5.1) that illustrates how added capacity can worsen outcomes, and an abstraction of the Amsterdam metro network (Section 5.2) to study commuter adaptation under different network phases. We conclude with a description of the experimental setup used to evaluate learning dynamics (Section 5.3).
@@ 5. Environments
Table 1. Overview of equilibrium results, without learning and $100$ agents starting at each source. We report total cost at the Nash Equilibrium (NE), Social Optimum (SO), Price of Anarchy (PoA, Eq. 7), average Source Disparity (SD, Eq. 11).
Figure 2. Top: The Amsterdam Metro Environment with key stations and routes for two commuter sources (West and East) converging at Central (See the Appendix for an overlay on the actual Metro Network Map). Three network phases are shown: Phase A (pre-2018, no North–South line), Phase B (current, with North–South line), and Phase C (future expansion connecting West and East via Amstel). Edge labels indicate free-flow travel times. Bottom: Arrow sizes indicate flows under the Nash Equilibrium, illustrating route choices and congestion distribution.
5.1. A Two-Source Braess’s Paradox Game
The Braess’s paradox demonstrates that adding capacity to a network can paradoxically increase congestion braess_paradox. Although often illustrated using stylized networks, it is neither rare nor anomalous under realistic assumptions steinberg1983prevalence, and has been observed empirically in cities such as New York, Boston, and Seoul youn2008price; baker2009removing.
To account for urban segregation effects and the observation that different demographic groups can systematically start commuting routes from different locations, we consider an extended Braess network with two source nodes and a single destination, illustrated in Figure 1. Two configurations are studied: (1) without the fast lane, where both sources have two strategies to reach the destination; and (2) with a zero-cost link from node $C$ to $D$, giving each source a third strategy.
@@ 5.1. A Two-Source Braess’s Paradox Game
The latency functions $f_{e}(x;K_{e})$ for edges $e\in E$ follow the general framework we defined in Section 3. Specifically: for edges $S_{1}\to C$ and $C\to B$, the latency is $f_{e}(x;K_{e})=x/N_{1}$; for edges $S_{2}\to D$ and $D\to B$, the latency is $f_{e}(x;K_{e})=x/N_{2}$; for edges $S_{1}\to D$ and $S_{2}\to C$, the latency is constant: $f_{e}(x;K_{e})=1$. When the additional fast lane $C\to D$ is introduced, it has constant latency $f_{e}(x;K_{e})=0$, mimicking the original Braess network.
@@
The game has non-decreasing latency functions, and thus, by Rosenthal’s theorem rosenthal1973network, admit at least one pure-strategy Nash equilibrium. We provide the detailed construction of the potential functions and full proof of equilibrium existence in the Appendix.
As shown in Table 1, in the baseline network, players from $S_{1}$ and $S_{2}$ converge to a socially optimal state. The intervention introduces a third strategy (Cross) for all players, shifting the Nash equilibrium so that all choose the new path, thereby increasing total social cost and the price of anarchy.
5.2. The Amsterdam Metro Network Game
To investigate how commuters adapt to changes in real-world transportation systems, we abstract the Amsterdam Metro Network into a routing game. Two origin groups of players start at $W$ and $E$, converging toward $C$ via intermediate hubs $S$ and $A$. Let $\mathcal{N}_{W}$ and $\mathcal{N}_{E}$ denote the player sets with $N_{W}$ and $N_{E}$ players, respectively, with $N=N_{W}+N_{E}$. The network structure is depicted in Figure 2 (top). See the Appendix (Figure 2) for details on how the real metro network is translated into the network we consider in this section.

### v2

@@ Abstract
Keywords Congestion Games $\cdot$ Reinforcement Learning $\cdot$ Multi-Agent Systems $\cdot$ Fairness $\cdot$ Transport Planning
1 Introduction
Expanding transport networks is often seen as an unambiguous good: more roads, more transport lines, more options means better access and faster travel. Yet, in practice, such interventions can backfire. Sometimes, adding infrastructure slows traffic—as showcased in Braess’s Paradox [10]—and other times, it disproportionately benefits already advantaged areas, widening inequalities [31].
@@ 1 Introduction
These challenges expose a critical blind spot in how we model and evaluate urban mobility. Classical game theoretic models assume homogeneous, rational agents with perfect information. But real commuters behave differently: they experiment and adapt. And although research that investigates this type of adaptation are increasing [11, 44, 6], they have not yet been analyzed through the lens of fairness, especially when agents differ in their capacity to learn. In this paper, we use multi-agent reinforcement learning on congestion games to investigate how exploration-driven learning shapes the distributional outcomes of transport expansions.
@@
A well-known example of unintended outcomes is the Braess’s Paradox, which shows that adding a new route can paradoxically worsen congestion. Initially a theoretical result, the paradox has since been observed in real-world cities like New York, Boston, and Seoul [50, 3]. Such paradoxes illustrate the power of game-theoretic models, which capture mobility as a strategic interaction among selfish agents [43, 51].
@@
Yet these models rest on the assumption of perfectly rational, fully informed commuters. In practice, travellers adapt through trial and error: experimenting with routes, observing outcomes, and adjusting their choices over time. This process closely resembles Reinforcement Learning (RL), in which agents balance exploration (trying new strategies) with exploitation (relying on known ones) [47]. Empirical studies have shown that even simple RL mechanisms can successfully reproduce observed patterns in commuter behaviour [44, 6]. Furthermore, modeling commuter behaviour through learning dynamics has shown deviations from the expected outcomes of full-information strategic agents — agents do not necessarily converge to the Nash Equilibrium [49, 11]. As such, introducing learning dynamics into traffic models not only increases realism but can also reshape expectations about system performance. Moreover, it enables the analysis throughout the entire adaptive process, rather than restricting attention solely to equilibria. We extend this line of research by tackling the fundamental question: How do differences in learning affect fairness and efficiency in congestion games?
@@
Our question is further motivated by empirical research suggesting that wealthier individuals showcase a higher risk tolerance, and are generally more exploratory in their behaviours [22, 9]. For instance, [22] find that residents of low-income neighborhoods are significantly less exploratory in their activities, effectively living within constrained environments. Psychological studies indicate that individuals from high-resource background demonstrate a higher tolerance for uncertainty and risk [9]. These findings imply that exploration and learning rates can act as a proxies for resource capacity: individuals with greater resources are better equipped to search for and evaluate alternatives, while those with fewer resources are constrained by limitations on time, mobility, and finances. Such heterogeneity can exacerbate inequalities, yet it remains underexplored.
@@ 2 Related Work
2.1 Reinforcement Learning in Congestion Games
@@
Deviations from the standard rational commuter behaviour have been studied under models such as prospect theory, which consider uncertainty in travelers’ decision-making [17, 16]. Recent papers revealed non-Nash convergence and even chaotic behaviour. For instance, online learning algorithms such as Multiplicative Weights Update [15] and Follow-the-Regularized-Leader [7] demonstrate that deviations from Nash equilibrium occur when considering learning agents. Learning automata have also been used [21].
@@
Reinforcement Learning is increasingly used to model travel choice behaviour in multiple congestion games [2]. Examples include studies on how commuters adapt over time through learning dynamics [13, 6], the impact of memory [48], as well as the impact of tolls in congestion games [37].
@@
Analyzing the effects of learning in congestion games through Reinforcement Learning, and in particular, multi-agent Q-learning hasn’t been explored thoroughly. Some efforts have achieved this in a parallel road network, showing that the total social cost varies with different exploration rates [29]. In particular, they show that with sufficiently small exploration, the system will converge to the inefficient equilibrium. Others investigate the effect of learning on a system’s convergence between Nash equilibrium and the social optimum [11]. However, while they assume a homogeneous population with a uniform learning rates, in reality, commuters can show heterogeneity in learning. Here, we build upon this work and extend this by modeling two distinct populations with varying learning rates, allowing for a richer analysis of strategic behaviour and its implications for equilibrium outcomes, on both a stylilzed and a real-world network.
2.2 Fairness in Congestion Games
@@
The challenge of avoiding increased costs after network expansions, as showcased by the well-known Braess’s Paradox, has been extensively studied assuming selfish agents, with efforts focused on network design [43, 26, 39].
@@
Our work contributes to this literature by examining fairness considerations in congestion games, an aspect previously explored primarily in centrally coordinated settings. Prior research has explored fairness through various mechanisms. One line of work analyzed the use of taxation to balance fairness and efficiency [24], while another investigated artificial currency–based pricing schemes as a means to promote equitable outcomes [36]. A karma-based framework for achieving fairness was proposed in related studies [14]. Further, the Double Braess’s Paradox was introduced in a multi-resource setting with two sources and two destinations, where action-restriction mechanisms were evaluated as a fairness intervention [35]. Additional work on weighted congestion games has examined mechanisms designed to reduce disparities in agent outcomes [25].
@@
Using agent-based models, prior work has examined how individual behaviors affect fairness and efficiency in network settings. One study analyzed a five-edge Braess’s paradox network, evaluating the impact of various micro-level parameters on system performance [5]. Another line of research investigated fairness in travel times between two routes through the lens of learning dynamics, assuming homogeneous exploration and learning rates [29].
@@
3 Preliminaries
3.1 Congestion Games
We consider standard congestion games [43]. Let $G=(V,E)$ be a directed graph, where $V$ is the set of nodes and $E$ the set of edges. Let $\mathcal{N}$ be the set of players with $N=|\mathcal{N}|$, where each player $i\in\mathcal{N}$ has a source node $s_{i}\in V$ and a common destination $d\in V$.
@@ 3.1 Congestion Games
  • Continuity and Monotonicity: $f_{e}$ is continuous and nondecreasing [42].
  • Polynomial-time computation: $f_{e}(x)$ can be computed in polynomial time [25].
@@ 3.2 Nash Equilibrium
Since congestion games are potential games [34, 33], at least one pure-strategy NE exists [40].
3.3 Social Optimum, Price of Anarchy, and Price of Learning
@@
The Price of Anarchy (PoA) measures inefficiency of NE relative to the social optimum [41]:
@@
where $a^{\star}$ is the socially optimal profile (Equation˜6).
@@
3.4 Fairness
We assess fairness through the notion of source disparity, defined as the difference in average cost between groups of players originating from different sources. Our definition is inspired by fairness (or justice) concepts in urban planning, where location is central: geographic position determines accessibility to amenities and thus shapes opportunities [31]. Our metric requires that no source node is systematically disadvantaged in the costs incurred to reach a destination. From the perspective of artificial intelligence fairness, this corresponds to the group fairness criterion of independence, which demands that outcomes—here, travel costs—remain independent of an agent’s point of origin [12].
@@ 3.4 Fairness
4 Learning Dynamics
In practice, humans deviate from exact Nash equilibrium strategies, as they lack complete information about which strategies yield specific payoffs [8, 28]. Instead of playing optimally from the outset, individuals tend to explore various options and gradually learn to favor those that maximize their satisfaction. This adaptive behavior has been studied in congestion games, where outcomes frequently diverge from Nash equilibrium predictions [11]. For this reason, we argue that it is essential to study the process itself, rather than assuming that commuters will ultimately converge. To achieve this, we use a framework in which agents repeatedly play the game and adapt their strategies through Reinforcement Learning (RL).
Importantly, we employ RL not as a tool for optimisation, but to simulate the adaptive behaviours of agents interacting within complex environments. Our perspective aligns with the descriptive agenda in multi-agent learning, which emphasizes modeling learning processes that mirror how real-world agents, such as humans, adapt in social contexts [45]. Importantly, this focus is not limited to human learning. The same framework extends to RL-based systems themselves. For example, as RL-powered recommender systems become increasingly prevalent [1], it is likely that real-time adaptive systems will also be deployed in navigation applications. Understanding their learning dynamics is therefore equally crucial.
@@ 4.1 Q-Learning Framework
with $a_{i}(t)$ the strategy chosen at time $t$ and $a_{t}$ the realized profile. The negative sign ensures Q-learning maximizes the negative latency (minimizes actual latency). We assume the game is stateless and rewards are observed immediately (as in, e.g., [18, 11]). Agents starting in different sources may have different learning rates.
@@
Convergence At each step, agents collectively contribute to congestion. Each agent estimates the expected cost of selecting a given strategy. With continued learning, Q-values converge to estimates reflecting expected cost. However, these do not necessarily correspond to a Nash Equilibrium (NE); convergence is shaped by factors such as the learning rate, exploration rate, and the environment’s structural properties [11]. Unlike traditional RL aimed at maximizing a reward, our focus here is on describing and understanding emergent phenomena in a population of learning agents.
Figure 1: A Braess’s Paradox game with two sources ($S1$ & $S2$) connecting to a single destination node $B$. We examine two phases: before and after the network extension $I_{CD}$ (red link).
5 Environments
We introduce two routing games: a stylized two-source extension of Braess’s paradox (Section˜5.1) that illustrates how added capacity can worsen outcomes, and an abstraction of the Amsterdam metro network (Section˜5.2) to study commuter adaptation under different network phases. We conclude with a description of the experimental setup used to evaluate learning dynamics (Section˜5.3).
@@ 5 Environments
Table 1: Overview of equilibrium results, without learning and $100$ agents starting at each source. We report total cost at the Nash Equilibrium (NE), Social Optimum (SO), Price of Anarchy (PoA, Eq. 7), average Source Disparity (SD, Eq. 11).
Figure 2: Top: The Amsterdam Metro Environment with key stations and routes for two commuter sources (West and East) converging at Central (See the Appendix for an overlay on the actual Metro Network Map). Three network phases are shown: Phase A (pre-2018, no North–South line), Phase B (current, with North–South line), and Phase C (future expansion connecting West and East via Amstel). Edge labels indicate free-flow travel times. Bottom: Arrow sizes indicate flows under the Nash Equilibrium, illustrating route choices and congestion distribution.
5.1 A Two-Source Braess’s Paradox Game
The Braess’s paradox demonstrates that adding capacity to a network can paradoxically increase congestion [10]. Although often illustrated using stylized networks, it is neither rare nor anomalous under realistic assumptions [46], and has been observed empirically in cities such as New York, Boston, and Seoul [50, 3].
To account for urban segregation effects and the observation that different demographic groups can systematically start commuting routes from different locations, we consider an extended Braess network with two source nodes and a single destination, illustrated in Figure˜1. Two configurations are studied: (1) without the fast lane, where both sources have two strategies to reach the destination; and (2) with a zero-cost link from node $C$ to $D$, giving each source a third strategy.
@@ 5.1 A Two-Source Braess’s Paradox Game
The latency functions $f_{e}(x;K_{e})$ for edges $e\in E$ follow the general framework we defined in Section˜3. Specifically: for edges $S_{1}\to C$ and $C\to B$, the latency is $f_{e}(x;K_{e})=x/N_{1}$; for edges $S_{2}\to D$ and $D\to B$, the latency is $f_{e}(x;K_{e})=x/N_{2}$; for edges $S_{1}\to D$ and $S_{2}\to C$, the latency is constant: $f_{e}(x;K_{e})=1$. When the additional fast lane $C\to D$ is introduced, it has constant latency $f_{e}(x;K_{e})=0$, mimicking the original Braess network.
@@
The game has non-decreasing latency functions, and thus, by Rosenthal’s theorem [40], admit at least one pure-strategy Nash equilibrium. We provide the detailed construction of the potential functions and full proof of equilibrium existence in the Appendix.
As shown in Table˜1, in the baseline network, players from $S_{1}$ and $S_{2}$ converge to a socially optimal state. The intervention introduces a third strategy (Cross) for all players, shifting the Nash equilibrium so that all choose the new path, thereby increasing total social cost and the price of anarchy.
5.2 The Amsterdam Metro Network Game
To investigate how commuters adapt to changes in real-world transportation systems, we abstract the Amsterdam Metro Network into a routing game. Two origin groups of players start at $W$ and $E$, converging toward $C$ via intermediate hubs $S$ and $A$. Let $\mathcal{N}_{W}$ and $\mathcal{N}_{E}$ denote the player sets with $N_{W}$ and $N_{E}$ players, respectively, with $N=N_{W}+N_{E}$. The network structure is depicted in Figure˜2 (top). See the Appendix (Figure 2) for details on how the real metro network is translated into the network we consider in this section.